# Traceable Cross-Source RAG for Chinese Tibetan Medicine Question Answering

### v1

@@ 1 Introduction
Large Language Models (LLMs) are increasingly used for information seeking and question answering. However, in high stakes domains, fluent responses are not enough. LLMs can hallucinate facts and fabricate evidence, which harms trust and safety (huang2025hallucination_survey, ; sahoo2024hallucination_survey, ; yu2024auepora, ). Retrieval Augmented Generation (RAG) mitigates this issue by grounding generation on external corpora. RAG has shown clear gains in clinical tasks when retrieval uses curated resources (zakka2024almanac, ; amugongo2025rag_healthcare_review, ). Yet recent analyses and benchmarks show that RAG remains fragile. Errors come from retrieval failures, noisy context, and unfaithful use of evidence (xiong-etal-2024-benchmarking, ; gao-etal-2023-enabling, ; es-etal-2024-ragas, ; yan2024crag, ; asai2024selfrag, ).
@@
Traditional medicine question answering has stronger requirements on provenance. In Chinese settings, Traditional Chinese Medicine (TCM) and Tibetan Medicine (TM) knowledge is scattered across heterogeneous sources. These sources include classical canons, modern clinical papers, and encyclopedic summaries. They differ in writing style, term usage, and authority. Recent work has built TCM and TM question answering systems with knowledge graphs (KGs) or LLM-based assistants (li2024dcmed_kgqa, ; qin2025ragcpmf, ; qin2025tcmlcm, ; dong2021tibetan_kgqa, ; pei2024tibetan_materia_medica_kg, ). In parallel, domain benchmarks and models for Chinese medical NLP and TCM evaluation are emerging (zhang2022cblue, ; wang2023huatuo, ; yu2024tcmd, ; wang2024domainrag, ). However, existing systems typically index a single source or merge all sources into one flat corpus. They rarely study how source heterogeneity affects retrieval choices, evidence fusion, and traceability. This leaves a clear gap for TM question answering in Chinese, where multi-source partitioned knowledge bases are common.
@@
This paper targets a practical but underexplored setting. We assume TM knowledge is split into separate knowledge bases by category, such as classics, clinical literature, and encyclopedia. Our goal is to make RAG answers more traceable, reduce hallucinations, and support cross-source verification. Two challenges are central. First, dense sources can dominate retrieval. In our data, encyclopedia passages are short and information dense, which makes retrievers prefer them. But for many clinical or doctrine sensitive questions, encyclopedic summaries are not the best evidence. This mismatch can lead to plausible but poorly grounded answers. This problem is related to resource selection in federated search, but the objective here also involves authority and provenance (wang2024feb4rag, ; wang2024resllm, ).
@@
Second, naive cross source fusion can reduce precision. A common approach is to concatenate retrieved passages from multiple sources and let the LLM decide. But long context models can be sensitive to passage order and position, which makes such fusion unstable (liu-etal-2024-lost-middle, ). Hybrid retrieval methods and KG-guided reranking are promising, but current work mainly evaluates general domains (yu2022kgfid, ; sarmah2024hybridrag, ). In TM, the problem is amplified by the semantic gap between classical terminology and modern clinical writing. As a result, simply adding more passages often increases noise and harms evidence attribution.
@@
We present a RAG framework designed for traceable TM question answering in Chinese under partitioned multi source knowledge bases. Our approach is evaluated with a dataset and metrics that explicitly measure correctness and traceability, beyond surface fluency (gao-etal-2023-enabling, ; es-etal-2024-ragas, ; xiong-etal-2024-benchmarking, ; yu2024auepora, ). Our contributions are:
@@ 2 Related Work
A central motivation for retrieval-augmented generation (RAG) is to improve factuality by grounding answers in external evidence, yet reliable attribution remains non-trivial. ALCE formalizes citation-aware generation and proposes automatic metrics to assess citation quality beyond answer fluency and correctness (gao-etal-2023-enabling, ). Complementarily, RAGAs provides a reference-free evaluation toolkit that decomposes RAG quality into retrieval and generation dimensions, enabling rapid iteration without expensive human labels (es-etal-2024-ragas, ). For domain settings, MIRAGE benchmarks medical RAG across many configurations and highlights that even strong pipelines can remain sensitive to component choices (xiong-etal-2024-benchmarking, ). These efforts motivate evaluation protocols that jointly consider answer correctness, evidence support, and citation reliability, which are critical for high-stakes knowledge domains.
@@
When multiple pieces of evidence are concatenated into a single prompt, large language models may underuse relevant information depending on its position. The “lost-in-the-middle” phenomenon shows that performance can degrade when key evidence is placed in the middle of long contexts, even for long-context models (liu-etal-2024-lost-middle, ). This observation is especially relevant for multi-source RAG, where naive concatenation or arbitrary ordering of retrieved passages can reduce precision. It motivates evidence organization strategies that are robust to ordering effects and can better support traceable, evidence-backed answers.
@@
Knowledge graphs (KGs) are often introduced into RAG to encode structured relations among entities and to improve evidence selection. HybridRAG combines vector retrieval with KG-based retrieval and reports improved retrieval and generation quality compared with using either signal alone (sarmah-etal-2024-hybridrag, ). In traditional medicine, recent work has started to build RAG-style question answering systems for Traditional Chinese Medicine (TCM), typically by constructing a single vectorized corpus from curated textbooks and classical materials (zhang-etal-2024-tcm-rag-llm, ). For Tibetan medicine, existing published systems more commonly focus on KG construction and KG-based question answering rather than multi-source RAG over partitioned corpora (pei2024tibetan_materia_medica_kg, ; dong2021tibetan_kgqa, ). Overall, prior work provides useful building blocks (citation evaluation, RAG assessment, long-context analysis, and KG-enhanced retrieval), but it leaves a gap in studying partitioned, heterogeneous traditional medicine knowledge bases where provenance, authority, and cross-library evidence alignment are first-order requirements.
@@ 4.2 Experimental Setup
We use openPangu-Embedded-7B as the answer generator for all methods chen2025pangu . This choice highlights a lightweight, deployment-friendly LLM in a domain RAG setting. To ensure a fair comparison, we keep the same generation prompt and citation format across all methods.

### v2

@@ 1 Introduction
Large Language Models (LLMs) are increasingly used for information seeking and question answering. However, in high stakes domains, fluent responses are not enough. LLMs can hallucinate facts and fabricate evidence, which harms trust and safety (Huang et al., 2025; Sahoo and others, 2024; Yu et al., 2024a). Retrieval Augmented Generation (RAG) mitigates this issue by grounding generation on external corpora. RAG has shown clear gains in clinical tasks when retrieval uses curated resources (Zakka et al., 2024; Amugongo and others, 2025). Yet recent analyses and benchmarks show that RAG remains fragile. Errors come from retrieval failures, noisy context, and unfaithful use of evidence (Xiong et al., 2024; Gao et al., 2023; Es et al., 2024; Yan et al., 2024; Asai et al., 2024).
@@
Traditional medicine question answering has stronger requirements on provenance. In Chinese settings, Traditional Chinese Medicine (TCM) and Tibetan Medicine (TM) knowledge is scattered across heterogeneous sources. These sources include classical canons, modern clinical papers, and encyclopedic summaries. They differ in writing style, term usage, and authority. Recent work has built TCM and TM question answering systems with knowledge graphs (KGs) or LLM-based assistants (Li and others, 2024; Qin and others, 2025a, b; Dong et al., 2021; Pei et al., 2024). In parallel, domain benchmarks and models for Chinese medical NLP and TCM evaluation are emerging (Zhang et al., 2022; Wang et al., 2023; Yu et al., 2024b; Wang and others, 2024). However, existing systems typically index a single source or merge all sources into one flat corpus. They rarely study how source heterogeneity affects retrieval choices, evidence fusion, and traceability. This leaves a clear gap for TM question answering in Chinese, where multi-source partitioned knowledge bases are common.
@@
This paper targets a practical but underexplored setting. We assume TM knowledge is split into separate knowledge bases by category, such as classics, clinical literature, and encyclopedia. Our goal is to make RAG answers more traceable, reduce hallucinations, and support cross-source verification. Two challenges are central. First, dense sources can dominate retrieval. In our data, encyclopedia passages are short and information dense, which makes retrievers prefer them. But for many clinical or doctrine sensitive questions, encyclopedic summaries are not the best evidence. This mismatch can lead to plausible but poorly grounded answers. This problem is related to resource selection in federated search, but the objective here also involves authority and provenance (Wang et al., 2024a, b).
@@
Second, naive cross source fusion can reduce precision. A common approach is to concatenate retrieved passages from multiple sources and let the LLM decide. But long context models can be sensitive to passage order and position, which makes such fusion unstable (Liu et al., 2024). Hybrid retrieval methods and KG-guided reranking are promising, but current work mainly evaluates general domains (Yu et al., 2022; Sarmah and others, 2024). In TM, the problem is amplified by the semantic gap between classical terminology and modern clinical writing. As a result, simply adding more passages often increases noise and harms evidence attribution.
@@
We present a RAG framework designed for traceable TM question answering in Chinese under partitioned multi source knowledge bases. Our approach is evaluated with a dataset and metrics that explicitly measure correctness and traceability, beyond surface fluency (Gao et al., 2023; Es et al., 2024; Xiong et al., 2024; Yu et al., 2024a). Our contributions are:
@@ 2 Related Work
A central motivation for retrieval-augmented generation (RAG) is to improve factuality by grounding answers in external evidence, yet reliable attribution remains non-trivial. ALCE formalizes citation-aware generation and proposes automatic metrics to assess citation quality beyond answer fluency and correctness (Gao et al., 2023). Complementarily, RAGAs provides a reference-free evaluation toolkit that decomposes RAG quality into retrieval and generation dimensions, enabling rapid iteration without expensive human labels (Es et al., 2024). For domain settings, MIRAGE benchmarks medical RAG across many configurations and highlights that even strong pipelines can remain sensitive to component choices (Xiong et al., 2024). These efforts motivate evaluation protocols that jointly consider answer correctness, evidence support, and citation reliability, which are critical for high-stakes knowledge domains.
@@
When multiple pieces of evidence are concatenated into a single prompt, large language models may underuse relevant information depending on its position. The “lost-in-the-middle” phenomenon shows that performance can degrade when key evidence is placed in the middle of long contexts, even for long-context models (Liu et al., 2024). This observation is especially relevant for multi-source RAG, where naive concatenation or arbitrary ordering of retrieved passages can reduce precision. It motivates evidence organization strategies that are robust to ordering effects and can better support traceable, evidence-backed answers.
@@
Knowledge graphs (KGs) are often introduced into RAG to encode structured relations among entities and to improve evidence selection. HybridRAG combines vector retrieval with KG-based retrieval and reports improved retrieval and generation quality compared with using either signal alone (Sarmah et al., 2024). In traditional medicine, recent work has started to build RAG-style question answering systems for Traditional Chinese Medicine (TCM), typically by constructing a single vectorized corpus from curated textbooks and classical materials (Zhang et al., 2024). For Tibetan medicine, existing published systems more commonly focus on KG construction and KG-based question answering rather than multi-source RAG over partitioned corpora (Pei et al., 2024; Dong et al., 2021). Overall, prior work provides useful building blocks (citation evaluation, RAG assessment, long-context analysis, and KG-enhanced retrieval), but it leaves a gap in studying partitioned, heterogeneous traditional medicine knowledge bases where provenance, authority, and cross-library evidence alignment are first-order requirements.
@@ 4.2 Experimental Setup
We use openPangu-Embedded-7B as the answer generator for all methods Chen et al. (2025). This choice highlights a lightweight, deployment-friendly LLM in a domain RAG setting. To ensure a fair comparison, we keep the same generation prompt and citation format across all methods.